OpenStudy (joie):

what is the distance of 0 to -6 and midpoint 1 to 5..i dont know how to solve it. show work please..

OpenStudy (anonymous):

distance from 0 to -6 = 0-(-6)=6 midpoint= (1+5)/2=6/2= 3

OpenStudy (anonymous):

did u get it joie?

OpenStudy (anonymous):

this is a example, If you have two points (x1, y1) and (x2, y2): (1,1) and (3,2): then you enter: Point 1: x1 = 1, y1 = 1 and Point 2: x2 = 3 y2 = 2. Click CALCULATE and slope is m = 0.5, line equation is y = 0.5x + 0.5, Distance = 2.2360679, and Midpoint = (2, 1.5). Enter your (x1, y1), (x2,y2) in the calculator and press Calculate.

OpenStudy (joie):

mark : yes...:D

OpenStudy (joie):

numbers:thank you for your example..